An ideal gas has molar heat capacity Cv at constant volume. The gas undergoes the process T=To (1-av) where a, To is constant. Then find the molar heat capacity of the gas.

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
C=Cv+(P*dV) /(ndT) C=Cv-R[R-(1-$V)]/[$V]$= alpha
Hope it helps
 
Regards
Arun (askIITians forum expert)